Step I Genetics probability question --- Math idiot O_o

Started by Mkandil
This forum made possible through the generous support of SDN members, donors, and sponsors. Thank you.
Get help with your application

Use all the free resources available to you from SDN: articles, guides, expert advising, forums discussions, and school research.

Mkandil

Full Member
10+ Year Member
Advertisement - Members don't see this ad
So I came across this question and tried to solve taking different approaches and it just doesn't compute. I know I'm an idiot in math and I need a second brain to help me solve this

A man and a woman are known to have an autosomal dominant disorders. They are both affected heterozygotes. The disorder has 80% penetrance. what are the chances that they will have a normal kid?

the correct answer is 40%

heterozygots of autosomal dominant diseases have 75% of passing the disease, multiplied by the 80% penetrance= 60% chances of having an affected kid. So the chances of having a normal kid = 100% - 60%= 40%

Now, trying different calculation
heterozygots of autosomal dominant disease have 25% of having a healthy kid, multiplied by the 20% non penetrance = 5 % !!!!

please tell me where I went wrong